OpenStudy (anonymous):

Plzzz Help? 9(d+5)=3(d-3)+6d a. empty set b.-3 c.all numbers d.-54

OpenStudy (anonymous):

i thought it was -54

OpenStudy (anonymous):

@Frvntic what are you tyrying to figure out

OpenStudy (anonymous):

im just tryna solve the equation

OpenStudy (anonymous):

Simplifying 9(d + 5) = 3(d + -3) + 6d Reorder the terms: 9(5 + d) = 3(d + -3) + 6d (5 * 9 + d * 9) = 3(d + -3) + 6d (45 + 9d) = 3(d + -3) + 6d Reorder the terms: 45 + 9d = 3(-3 + d) + 6d 45 + 9d = (-3 * 3 + d * 3) + 6d 45 + 9d = (-9 + 3d) + 6d Combine like terms: 3d + 6d = 9d 45 + 9d = -9 + 9d Add '-9d' to each side of the equation. 45 + 9d + -9d = -9 + 9d + -9d Combine like terms: 9d + -9d = 0 45 + 0 = -9 + 9d + -9d 45 = -9 + 9d + -9d Combine like terms: 9d + -9d = 0 45 = -9 + 0 45 = -9 Solving 45 = -9

OpenStudy (anonymous):

so its a no solution.

OpenStudy (anonymous):

i was doing that lol looks like you beat me to it

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

wait no what it he doing with the d \

OpenStudy (anonymous):

thnx

OpenStudy (anonymous):

the d tshould remain as 1

OpenStudy (anonymous):

any number next to a letter means the letter is 1

OpenStudy (anonymous):

yeah but they got canceled out